The sum of the measures of the angles of a quadrilateral is 360 degrees. Suppose you are told tha tthe angle measures of a certain quadrilateral are consecutive multiples of 10 (such as 10,20,30.) show that this is impossible.
Answer by Alan3354(31532)

x + x+10 + x+20 + x+30 = 360
4x + 60 = 360
x = 75
--> 75, 85, 95 & 105 not multiples of 10
--------------
The avg has to be 360/4 = 90
4 numbers with an average of 90 are not consecutive multiples of 10